OpenStudy (anonymous):

After a discount of 20%, the price of fitness equipment, including tax, was $73.53. The following equation can be used to find x, the price of the equipment without the discount. x - 0.2x + 0.07(x - 0.2x) = 73.53 What is the price of the fitness equipment without the discount?

OpenStudy (anonymous):

let before discount the original price be x so 20% of x =73.53 \[\frac{20}{100} \times x = 73.53 \rightarrow x = \frac{73.53 \times 100}{20} =73.53 \times 5 =367.65\] Thus, before discount the original price = 367.65 dollars Now you substitute x= 367.65 in the given eq and you will find your required Ans.